OpenStudy (solomonzelman):

Please help -- very basic. I'm stuck w/ some q abt recursive formula for the sequence:

OpenStudy (solomonzelman):

How do I write a recursive formula for \[\sqrt{p+q\sqrt{p+q\sqrt{p+q...}}}\](just for clarity the terms are:\[a_1=\sqrt{p+q}\]\[a_2=\sqrt{p+q\sqrt{p+q}}\]\[a_3=\sqrt{p+q\sqrt{p+q\sqrt{p+q}}}\]and I continue going like this, multiplying the very last q in \(a_{n-1}\), times \(\sqrt{p+q}\) to get the \(a_n\).

OpenStudy (rational):

try \[a_{n+1}=\sqrt{p+\sqrt{a_n}}; ~~a_1={\sqrt{p+q}}\]

OpenStudy (solomonzelman):

oh, that makes perfect sense.... I wasn't thinking about that form of it. tnx

OpenStudy (solomonzelman):

So, \[a_1=p+q\]\[a_2=p+\frac{q}{p+q}\]\[a_3=p+\frac{q}{p+\frac{q}{p+q}}\] and here the Formula is\[a_3=p+\frac{q}{a_n};~~a_1=p+q\]

OpenStudy (solomonzelman):

I mean not a_3 in te last line, but a_(n+1)

OpenStudy (solomonzelman):

\[a_{n+1}=p+\frac{q}{a_n};~~a_1=p+q\]

OpenStudy (rational):

looks good!

OpenStudy (solomonzelman):

yeah, I guess I wasn't taught properly... now I got another concept in my closet. thanks again, and best of luck !

OpenStudy (rational):

you're doing good, yw!